# Enhancing Multimodal Large Language Models with Vision Detection Models: An Empirical Study

### v2

@@ 4.5 Overview of More Experiments
In Appendix C.1, we fine-tune LLaVA-1.5 without the infusion of detection information and discover that the exceptional performance of FTBI models is primarily ascribed to the infused detection information, rather than the additional fine-tuning. In Appendix C.2, we assess the performance of FTBI models without infusing any detection information, affirming that the FTBI strategy proficiently harnesses ViT features. In Appendix C.3, we show the model’s performance on solely leveraging object detection information and OCR information separately. In Appendix C.4, we assess our models’ inference efficiency, showing that even with the detection infusion, our approach remains efficient with potential for further efficiency gains through modular replacement and text compression strategies.
Model Architecture Rationale.
In Appendix D.1, we discuss how does LLaVA-1.5 represents the majority of advanced MLLMs, supported by their architecture alignment and more empirical results on another MLLM, Qwen-VL. In Appendix D.2, we demonstrate the representativeness of DINO and PaddleOCRv2 to other models, thanks to the proposed flexible format that allows us to process the outputs of any detection models into detection information. Besides, we replace DINO with another detector, YOLOv8 [20], yielding similar experimental findings. In Appendix D.3, we show that using text as a bridge to feed detection information into LLMs generally gain superior results without the need for additional data, compared with feeding methods based on MLP or cross-attention training.
@@
In Appendix E.1, we elaborate on the motivations and modifications behind the GQA from the original GQA benchmark. In Appendix E.2, we present detailed MME benchmark scores. In Appendix E.3, we evaluate our model’s ability to ground specific linguistic phenomena with the VALSE benchmark [41]. In Appendix E.4, we explore the variants of calculation of aggregated score $\bar{s}_{norm}$, based on importance weights of benchmarks.
4.6 Limitation and Future Work
Refinement of Detection Models.
The success of our model relies heavily on the accuracy and reliability of the incorporated detection models. However, existing detection models are not flawless, as they still make mistakes when encountering challenging issues. For instance, our model’s improvement on the TextVQA benchmark is not very significant, primarily due to TextVQA being a scene-text dataset, and PaddleOCRv2 performs poorly on such data. Fortunately, both object detection and OCR models are still rapidly advancing [62]. It’s believed that highly accurate and fast models will emerge in the near future. Therefore, future work could explore the development and infusion of more advanced detection models to further improve the MLLMs.
Extended Token Sequences.
The infusion of detection information into token sequences extends the typical input length for the backbone LLM, potentially impacting inference speed negatively. Nonetheless, our statistical investigation reveals that the detection data, thanks to compressed prompts, mostly remain succinct. Furthermore, advancements in AI hardware, particularly the rapid evolution of AI chips, coupled with ongoing research efforts to expand the contextual capacity of LLMs [17], are progressively alleviating the imposed constraints. Future work could upgrade our MLLMs to support LLMs with longer contextual windows, further enhancing the performance.
@@ 5 Conclusion
In this paper, we systematically conduct substantial experiments to find effective strategies for incorporating SOTA object detection and OCR models into MLLMs. We validate the effectiveness of leveraging text as an intermediary to infuse detection information. After thorough investigation, we determine that fine-tuning the original MLLM for an additional epoch, along with the simultaneous infusion of detection information, proves to be the most effective approach compared to the training-free strategy and retraining strategy. Moreover, we replace the detection model from DINO to GroundingDINO and observe that the updated model continues to function properly and retains the training effect. This highlights the modular structure of our model, demonstrating its capability to stay abreast of evolving object detection technologies and gain sustained performance enhancement.
We offer guidance on how to infuse detection information into MLLMs. If people want to enhance the zero-initialized MLLMs, they can consider introducing detection modules throughout the training process to achieve effects similar to those of RBI models. To further improve the performance, they can initially train the model without detection information to develop robust non-detail-related capabilities. Then, they can introduce detection modules and fine-tune the model to develop visual grounding capabilities. If people aim to enhance pre-trained MLLMs, they can directly conduct additional fine-tuning on the model while infusing detection information simultaneously.
In a nutshell, we provide a series of progressive insights about the effective incorporation of vision detection models for MLLMs, and derive models that exhibit exceptional performance in various tasks such as counting, object localization, and text recognition. With this work, we hope it can benefit future MLLM research and development that approaches better understanding, interpreting and engaging with fine-grained multimodal content.
@@
  • Appendix A.1: we provide examples on which LLaVA-1.5-13B fails while our model FTBI-13B with detection information succeeds.
  • Appendix A.2: we provide examples of images along with their corresponding detection information, showing how the detection information is constructed.
@@
  • Appendix B.1: we conduct a statistical analysis on the length of detection information, showcasing the efficacy of our compression strategy.
@@
  • Appendix B.3: we show the different input resolutions of images across three branches: CLIP-ViT, DINO, and PaddleOCRv2.
@@
  • Appendix B.5: we list the training hyperparameters of our training strategies.
@@
  • Appendix C.1: we fine-tune LLaVA-1.5 without the infusion of detection information and get new models. The results indicate that the exceptional performance of FTBI models is primarily ascribed to the infused detection information, rather than the additional fine-tuning conducted on LLaVA-1.5.
  • Appendix C.2: we assess the performance of FTBI models without infusing any detection information, affirming that the FTBI fusion strategy proficiently harnesses the features extracted from ViT.
  • Appendix C.3: we show the performance of FTBI models exclusively infusing OCR information or object detection information, affirming that they can respectively enhance the performance of our models on relevant tasks.
  • Appendix C.4: we assess the inference efficiency of our models with detection information.
@@
  • Appendix D.1: we discuss how LLaVA-1.5 represents the majority of advanced MLLMs, and the results of LLaVA-1.5 can be extended to other MLLMs with similar structures. Additionally, we perform experiments on another MLLM, Qwen-VL, validating the versatility of our paper’s experimental findings.
  • Appendix D.2: we show how do DINO and PaddleOCRv2 represent other detection models in our experiments. Additionally, we perform experiments on another detection model, YOLO-v8, validating that the specific format we devise for processing detection information to reduce the importance of model selection.
  • Appendix D.3: we investigate other optional approaches to feed detection information into the LLM, comparing with our adopted text-based approach on the main page.
@@
  • Appendix E.1: we elaborate on the rationals to modify the GQA benchmark.
  • Appendix E.2: we present benchmark scores for RBI models and FTBI models on MME-Perception and MME-Cognition.
  • Appendix E.3: we evaluate our models’ ability to ground specific linguistic phenomena with the VALSE benchmark.
  • Appendix E.4: we introduce weights to calculate the aggregated score based on the importance of each benchmark.
@@ Appendix A More Demonstrative Examples
A.1 Examples on which LLaVA-1.5-13B fails while FTBI-13B with detection information succeeds.
Table 5 presents examples where LLaVA-1.5 provides incorrect responses, while our FTBI-13B delivers accurate answers. The showcased examples encompass scenarios related to object counting, object detection, and text recognition.
Figure 5: Examples on which LLaVA-1.5-13B fails while our model with detection information (FTBI-13B) succeeds.
A.2 Examples of Detection Information
Figure 6: The composition of detection information. “INS”, “obj/cls” and “pos” indicate instruction, detected object/class name, and position text respectively.
Figure 7: Examples of detection information generated with DINO and PaddleOCRv2.
@@ Appendix B Implementation Details
B.1 Length of Detection Information
Since the textual descriptions of bounding box coordinates typically involve a lot of digits, their token sequences are often long. As introduced in Section 3.2, we devise strategies to succinctly represent the spatial information of detected objects and texts, mitigating the verbosity of bounding box descriptions. By focusing on central coordinates and consolidating objects within the same category, we maintain brevity and clarity in our model’s inputs.
@@ B.1 Length of Detection Information
The instruction-following dataset of LLaVA-1.5 [30] is a combination of several datasets that relate to various tasks. Among them, the LLaVA dataset [31] and ShareGPT dataset [7] comprise high-quality GPT-4 conversation data. VQAv2 [14] and GQA [18] present samples that require one word or a short phrase to answer visual questions. OKVQA [39] and A-OKVQA [44] are VQA datasets designed to expand the knowledge base of multimodal models through the incorporation of external prior knowledge. OCRVQA [40] is expressly tailored to enhance the text recognition capabilities of multimodal models. TextCaps [45] is an image captioning dataset, which presents samples in the form of concise one-sentence descriptions corresponding to images. RefCOCO [21] and VG [22] are object detection datasets designed to improve the object localization capabilities of multimodal models.
@@
Regarding the pretraining dataset of LLaVA-1.5, it is pertinent to note that this dataset predominantly consists of samples tailored for image captioning, thus inherently emphasizing image-level information. However, our detection information focuses more on fine-grained details, so we opt not to incorporate this dataset in our FTBI fusion strategy.
@@ B.3 Image Resolution
  • For DINO and Grounding DINO’s preprocessing, input images can have arbitrary aspect ratios, but we need to limit the length of the shortest side to at least 224 and the length of the longest side to be within 2048. The setting for the shortest side length is to prevent insufficient multi-scale features extracted by DINO’s image encoder module, ensuring an adequate number of anchor boxes. The setting for the longest side length is to reduce additional memory usage, and this value can be set arbitrarily.
@@ B.4 Threshold Setting for Detection Models
We set certain thresholds for the detection models to reduce the acquisition of error information. Specifically, we set the threshold for DINO to 0.3 and only targets with confidence scores higher than this threshold are considered valid targets. For PaddleOCR, we set the detection box threshold to 0.6 and only detection boxes with confidence scores higher than this threshold are considered to contain text. For Grounding DINO, we set the detection box threshold to 0.35 and the text threshold to 0.25, and only targets meeting the requirements of both thresholds are considered valid targets.
@@ B.5 Training Hyperparameters
In Table 7, we show the training hyperparameters employed in our experiments. These hyperparameters are derived from Vicuna [7] and LLaVA-1.5 [30] and have proven to be effective in our experiments. In the table, the term “Pretrain-RBI” denotes the hyperparameters used during the pretraining phase for vision-language alignment in RBI fusion strategy. “Finetune-RBI” refers to the hyperparameters employed for the subsequent fine-tuning phase focusing on visual instruction tuning in RBI fusion strategy. Additionally, “Finetune-FTBI” designates the hyperparameters used during the fine-tuning process for FTBI fusion strategy.
@@
B.6 Time Consumption for Training
As for the experimental cost, on the 4xA100 (80G) platform, the pretraining process takes around 11 hours for 7B models and around 17 hours for 13B models. As for fine-tuning, it demands around 22 hours for 7B models and around 33 hours for 13B models.
Appendix C Further Experiments and Analysis on the FTBI Model
@@
Table 8: If we finetune LLaVA-1.5 without detection information, the performance will be inferior to the version with detection information. “-T w/o DI” stands for “training without detection information.”
@@
In FTBI fusion strategy, our models undergo an additional epoch of fine-tuning on LLaVA-1.5. In the current experiment, we train a different version of our FTBI models without the infusion of detection information during training. In this way, we can investigate whether the performance improvement of our FTBI models is attributable to the supplementary detection information or to the fine-tuning of an additional epoch.
As indicated in Table 8, the performances of the models without detection information are on par with LLaVA-1.5. Compared to FTBI models, these models exhibit inferior performance across almost all benchmarks. Consequently, the outstanding performance of our FTBI models is more attributed to the detection information we supplement, rather than that we fine-tune for an extra epoch on LLaVA-1.5.
C.2 Analysis of FTBI Models without Detection Information
Table 9: If we do not infuse detection information to FTBI-7B during inference, its performance will be on par with LLaVA-1.5-7B. “w/o DI” is an abbreviation for “without detection information.”
MLLM VQAv2 GQA* VQAT POPE MMEP MMEC MMB MMBCN MM-Vet SEED
LLaVA-1.5-7B 78.5 79.6 58.2 85.9 1510.7 355.7 64.3 58.3 30.5 58.6
RBI-7B w/o DI 76.4 74.8 56.6 85.5 1387.7 312.5 65.5 58.3 29.0 59.6
FTBI-7B w/o DI 78.0 $\uparrow$ 78.4 $\uparrow$ 57.1 $\uparrow$ 86.0 $\uparrow$ 1441.8 $\uparrow$ 303.6 $\downarrow$ 66.9 $\uparrow$ 59.7 $\uparrow$ 30.1 $\uparrow$ 60.6 $\uparrow$
FTBI-7B 79.0 80.1 59.8 88.9 1482.7 397.9 67.3 60.2 35.2 60.8
Table 10: If we do not infuse detection information to FTBI-13B during inference, its performance will be on par with LLaVA-1.5-13B. “w/o DI” is an abbreviation for “without detection information.”
[…]
LLaVA-1.5-13B 80.0 81.0 61.3 85.9 1531.3 295.4 67.7 63.6 35.4 61.6
RBI-13B w/o DI 77.3 76.0 58.2 83.4 1442.6 310.7 68.5 61.7 30.6 61.6
FTBI-13B w/o DI 79.4 $\uparrow$ 80.0 $\uparrow$ 60.0 $\uparrow$ 85.3 $\uparrow$ 1525.7 $\uparrow$ 320.0 $\uparrow$ 70.8 $\uparrow$ 64.8 $\uparrow$ 36.0 $\uparrow$ 61.7 $\uparrow$
FTBI-13B 80.3 81.8 61.8 88.8 1555.1 365.4 71.4 65.2 38.9 62.3
We assess the benchmark scores of FTBI models without infusing detection information to evaluate their capacities in leveraging ViT features. The findings delineated in Table 9 and Table 10 demonstrate that the efficacy of FTBI models without detection information aligns closely with that of LLaVA-1.5, and they outperform RBI models without detection information across all benchmarks. This implies that our third fusion strategy effectively empowers the model to assimilate and capitalize on information extracted by ViT.
C.3 Performance of FTBI Models Exclusively with OCR or Object Detection Information
Table 11: Performance of FTBI models only with OCR information.
@@ C.3 Performance of FTBI Models Exclusively with OCR or Object Detection Information
Table 12: Performance of FTBI models only with object detection information.
@@
As evident from Table 11 and Table 12, the inclusion of object detection information significantly boosts the scores of relevant benchmarks for object localization and object hallucination. Similarly, the incorporation of OCR information markedly improves the scores of benchmarks related to text recognition.
C.4 Inference Efficiency
We assess the time consumption of our FTBI-7B model by calculating its end-to-end inference time on the GQA and TextVQA datasets. When our model relies solely on object detection information during inference, DINO accounts for 38% of the total inference time. Additionally, when OCR information is exclusively infused, PaddleOCRv2 accounts for 25% of the total inference time. It is also noteworthy that, although the infusion of detection information incurs additional inference cost, as shown in Table 4, our model outperforms many state-of-the-art (SOTA) 13B or even larger multimodal large language models, which require considerably more resources. Moreover, thanks to the modularity and substitutability of our approach, a lighter and more efficient detection model could further improve the efficiency [49]. Additionally, since the embedded detection models are mutually independent, we can let them run independently on different devices, enabling parallel inference and further accelerating inference speed.
Regarding the proposed text compression strategy (Section 3.2), compared to using the original text output from detection models, the model with text compression achieves a 9% reduction in inference time when combined with object detection information, and a significant 58% reduction in inference time when combined with OCR information, verifying the effectiveness of the proposed strategy.

### v3

@@ 5 Overview of More Experiments
In Appendix C.1, we fine-tune LLaVA-1.5 without the infusion of detection information and discover that the exceptional performance of FTBI models is primarily ascribed to the infused detection information, rather than the additional fine-tuning. In Appendix C.2, we show the model’s performance on solely leveraging object detection information or OCR information.
@@
In Appendix E.1, we elaborate on the motivations and modifications behind the GQA. In Appendix E.2, we present detailed MME benchmark scores. In Appendix E.4, we evaluate our models’ ability to ground specific linguistic phenomena with the VALSE benchmark. In Appendix E.3, we evaluate the models on two DocumentVQA benchmarks, DocVQA and InfographicVQA.
@@ 6 Conclusion
In this paper, we systematically conduct experiments to compare the effects of different training strategies on the infusion of textual detection information into MLLMs. After thorough investigation, we determine that fine-tuning the original MLLM for an additional epoch, along with the simultaneous infusion of textual detection information, proves to be the most effective approach compared to the training-free strategy and the retraining strategy. Moreover, we replace the detection model deployed in the studied MLLM from a close-set detector to an open-set detector and observe that the updated fine-tuned model retains the training effect and achieve better performance than the updated training-free one. This indicates that the fine-tuned model, compared to the training-free model, can better stay abreast of evolving object detection technologies and achieve sustained performance enhancements.
In a nutshell, we provide a series of progressive insights about the effective infusion of textual detection information into MLLMs. We aim to inform researchers that when attempting to convert the outputs of vision detection models into textual information for assisting MLLMs, it can be beneficial to use a small amount of general VQA data for additional fine-tuning (potentially using the instruction-tuning data from the MLLM itself). This approach can yield models that perform better than those not subjected to training, allowing the models to have a more comprehensive understanding and utilization of the detection information. With this work, we hope it can benefit future MLLM research and development that approaches better understanding, interpreting and engaging with fine-grained multimodal content.
@@
  • Appendix A.1: we show examples on which LLaVA-1.5-13B fails while the model infused with textual detection information (FTBI-13B) succeeds.
  • Appendix A.2: we show examples of images and their corresponding textual detection information, illustrating how the textual detection information is constructed.
@@
  • Appendix B.1: we conduct a statistical analysis on the length of textual detection information, showcasing the efficacy of our compression strategy.
@@
  • Appendix B.3: we show the different input resolutions of three branches: CLIP-ViT, DINO, and PaddleOCRv2.
@@
  • Appendix B.5: we list the training hyperparameters.
@@
  • Appendix C.1: we fine-tune LLaVA-1.5 without the infusion of detection information and test the newly got models. The results indicate that the exceptional performance of FTBI models is primarily ascribed to the infused detection information, rather than the additional fine-tuning.
  • Appendix C.2: we show the performance of FTBI models exclusively infusing OCR information or object detection information, affirming that they can respectively enhance the performance of MLLMs on relevant tasks.
  • Appendix C.3: we assess the inference efficiency of the MLLM infused with textual detection information.
@@
  • Appendix D.1: we discuss how LLaVA-1.5 represents the majority of advanced MLLMs, and the results of LLaVA-1.5 can be extended to other MLLMs with similar structures. Additionally, we perform experiments on other MLLMs, Qwen-VL and LLaVA-NeXT, validating the versatility of our paper’s experimental findings.
  • Appendix D.2: we show how do DINO and PaddleOCRv2 represent other detection models in our experiments. Additionally, we perform experiments on another object detection model, YOLO-v8N, validating that the specific format we devise for processing textual detection information reduces the importance of model selection.
  • Appendix D.3: we conduct experiments based on YOLOv5N and YOLOv11L, and investigate the impact of detector accuracy on MLLM performance.
  • Appendix D.4: we remove the detection data from the instruction tuning dataset and repeat the FTBI experiment, aiming to investigate whether the model can still maintain good language comprehension capability.
  • Appendix D.5: we introduce LoRA modules to the visual encoder and repeat the retraining and fine-tuning experiments, obtaining results consistent with the conclusions presented on the main page
  • Appendix D.6: we conduct experiments based on Co-DETR-LVIS, which is capable of detecting 1.2K object categories, to explore the impact of a broader object detection scope on MLLM performance.
  • Appendix D.7: we discuss the main focus of our research, which is “deploying detection models to assist MLLMs’’, and highlight its distinction from related works that introduce special tokens to guide MLLMs. Furthermore, we elaborate on the motivation behind focusing our research on this paradigm.
@@
  • Appendix E.1: we provide further details about why we modify the GQA benchmark.
  • Appendix E.2: we present benchmark scores on MME-Perception and MME-Cognition.
  • Appendix E.3: we evaluate our models on two well-known DocumentVQA benchmarks, DocVQA and InfographicVQA.
  • Appendix E.4: we evaluate the models’ ability to ground specific linguistic phenomena with the VALSE benchmark, further confirming that the fine-tuning strategy is better than the training-free strategy.
@@ Appendix A More Demonstrative Examples
A.1 Examples on which LLaVA-1.5-13B fails while the model infused with textual detection information succeeds.
Table 5 presents examples where LLaVA-1.5 provides incorrect responses, while the FTBI-13B delivers accurate answers. The showcased examples encompass scenarios related to object counting, object detection, and text recognition.
Figure 5: Examples on which LLaVA-1.5-13B fails while the model infused with textual detection information (FTBI-13B) succeeds.
A.2 Examples of Textual Detection Information
Figure 6: The composition of textual detection information. “INS”, “obj/cls” and “pos” indicate instruction, detected object/class name, and position text respectively.
Figure 7: Examples of textual detection information generated with DINO and PaddleOCRv2.
@@ Appendix B Implementation Details
B.1 Length of Textual Detection Information
Since the textual descriptions of bounding box coordinates typically involve a lot of digits, their token sequences are often long. As introduced in Section 3, we devise strategies to succinctly represent the spatial information of detected objects and texts, mitigating the verbosity of bounding box descriptions. By focusing on central coordinates and consolidating objects within the same category, we maintain brevity and clarity in our model’s inputs.
@@ B.1 Length of Textual Detection Information
Dataset quality and diversity have been proven to be pivotal to the performance of MLLMs Chen et al. (2024a; b). The instruction-following dataset of LLaVA-1.5 (Liu et al., 2023a) is a meticulously curated combination of multi-task datasets, enhanced in both quality and diversity. Among them, the LLaVA dataset (Liu et al., 2023b) and ShareGPT dataset (Chiang et al., 2023) comprise high-quality GPT-4 conversation data. VQAv2 (Goyal et al., 2017) and GQA (Hudson & Manning, 2019) present samples that require one word or a short phrase to answer visual questions. OKVQA (Marino et al., 2019) and A-OKVQA (Schwenk et al., 2022) are VQA datasets designed to expand the knowledge base of multimodal models through the incorporation of external prior knowledge. OCRVQA (Mishra et al., 2019) is expressly tailored to enhance the text recognition capabilities of multimodal models. TextCaps (Sidorov et al., 2020) is an image captioning dataset, which presents samples in the form of concise one-sentence descriptions corresponding to images. RefCOCO (Kazemzadeh et al., 2014) and VG (Krishna et al., 2017) are object detection datasets designed to improve the object localization capabilities of multimodal models.
@@
Regarding the pretraining dataset of LLaVA-1.5, it is pertinent to note that this dataset predominantly consists of samples tailored for image captioning, thus inherently emphasizing image-level information. However, our detection information focuses more on fine-grained details, so we opt not to incorporate this dataset in our FTBI training strategy.
@@ B.3 Image Resolution
  • For DINO and Grounding DINO’s preprocessing, input images can have arbitrary aspect ratios. However, we need to limit the length of the shortest side to at least 224 and the length of the longest side to be within 2048. The setting for the shortest side length is to prevent insufficient multi-scale features extracted by DINO’s image encoder, ensuring an adequate number of anchor boxes. The setting for the longest side length is to reduce additional memory usage, and this value can be set arbitrarily.
@@ B.4 Threshold Setting for Detection Models
We set certain thresholds for the detection models to reduce the acquisition of error information. Specifically, we set the threshold for DINO to 0.3 and only targets with confidence scores higher than this threshold are considered valid targets. For PaddleOCR, we set the bounding box threshold to 0.6 and only bounding boxes with confidence scores higher than this threshold are considered to contain text. For Grounding DINO, we set the bounding box threshold to 0.35 and the text threshold to 0.25, and only targets meeting the requirements of both thresholds are considered valid targets.
@@ B.5 Training Hyperparameters
In Table 6, we show the training hyperparameters employed in our experiments. These hyperparameters are derived from Vicuna (Chiang et al., 2023) and LLaVA-1.5 (Liu et al., 2023a) and have proven to be effective. In the table, the term “Pretrain-RBI” denotes the hyperparameters used during the pre-training phase for vision-language alignment in RBI training strategy. “Finetune-RBI” refers to the hyperparameters employed for the subsequent fine-tuning phase focusing on visual instruction tuning in RBI training strategy. Additionally, “Finetune-FTBI” designates the hyperparameters used during the fine-tuning process for FTBI training strategy.
@@
B.6 Time Consumption and Memory Requirements
As for the training cost, on four NVIDIA A100 GPUs (80GB VRAM), the time consumption in terms of the original cost and with the detection information infusion is as follows:
  • For pretraining LLaVA-1.5-7B, the time increases from 6 hours to 11 hours.
  • For pretraining LLaVA-1.5-13B, the time increases from 11 hours to 17 hours.
  • For fine-tuning LLaVA-1.5-7B, the time increases from 16 hours to 22 hours.
  • For fine-tuning LLaVA-1.5-13B, the time increases from 26 hours to 33 hours.
Regarding the memory requirements, deploying detection models results in an additional GPU memory usage of up to 4GB in each GPU compared to not deploying detection models.
Appendix C Further Experiments and Analysis on the FTBI Models
@@
For the FTBI training strategy, the models undergo an additional epoch of fine-tuning based on LLaVA-1.5. In the current experiment, we will train a different version of FTBI models without the infusion of detection information during training. In this way, we can investigate whether the performance improvement of the FTBI models is attributable to the supplementary detection information or to the fine-tuning of an additional epoch.
Table 7: If we finetune LLaVA-1.5 without infusing textual detection information, the performance will be inferior to the version with detection information. “-T w/o DI” stands for “training without detection information.”
@@
As indicated in Table 7, the performance of the models fine-tuned without infusing detection information is on par with that of LLaVA-1.5. Compared to FTBI models, these models exhibit inferior performance across almost all benchmarks. Consequently, the outstanding performance of the FTBI models is more attributed to the textual detection information we supplement, rather than that we fine-tune for an extra epoch on LLaVA-1.5.
C.2 Performance of FTBI Models Exclusively with OCR or Object Detection Information
Table 8: Performance of FTBI models only infused with OCR information.
@@ C.2 Performance of FTBI Models Exclusively with OCR or Object Detection Information
Table 9: Performance of FTBI models only infused with object detection information.
@@
As evident from Table 8 and Table 9, the infusion of object detection information boosts the scores of relevant benchmarks for object localization and object hallucination. Similarly, the infusion of OCR information improves the scores of benchmarks related to text recognition.
C.3 Inference Efficiency
We assess the time consumption of the FTBI-7B model by calculating its end-to-end inference time with the GQA dataset and the TextVQA dataset. When the model relies solely on object detection information during inference, DINO accounts for 38% of the total inference time. Additionally, when OCR information is exclusively infused, PaddleOCRv2 accounts for 25% of the total inference time.
Thanks to the modularity of the studied MLLM architecture and the detection model replaceability enabled by the fine-tuning strategy, a lighter and more efficient detection model could further improve the efficiency (Wang et al., 2023b). Additionally, since the embedded detection models are mutually independent, we can let them run independently on different devices, enabling parallel inference and further accelerating inference speed.
Regarding the proposed text compression strategy (Section 3), we compare its performance with that of using the entire output from detection models (without selecting the first two values of coordinates). We find that the model with text compression achieves a 9% reduction in inference time when combined with object detection information, and a significant 58% reduction in inference time when combined with OCR information, verifying the effectiveness of the proposed text compression strategy.